# Discrete flavour and CP symmetries in light of JUNO and neutrino global fit

### v1

@@ Abstract
Working within the reference three-neutrino mixing framework, we confront the lepton mixing predictions derived using non-Abelian discrete flavour and CP symmetries with the firs JUNO data on the solar neutrino mixing parameters $\sin^{2}\theta_{12}$ and with the results of the latest global neutrino data analysis. We focus on symmetry breaking patterns for which the lepton PMNS mixing matrix depends only on one or two free real parameters. Performing a comprehensive statistical analyses in each of the considered cases, we report the best fit values, the $3\sigma$ C.L. allowed ranges and the $\chi^{2}$-distributions of the lepton mixing observables - the three mixing angles and the three CP-violation phases. We find that the JUNO measurements can disfavour or rule out a number of the mixing patterns associated with specific types of breakings of the discrete flavour and CP symmetries. The synergy of JUNO, DUNE and T2HK data can provide an exhaustive test of the considered approach to lepton mixing based on non-Abelian discrete lepton flavour symmetries combined with the CP symmetry.
@@ 1 Introduction
The neutrino oscillation data accumulated over many years of measurements following the discovery of neutrino oscillations [1, 2] revealed the existence of three-flavour neutrino mixing involving three light massive neutrinos (see, e.g., [3]). In this reference three-neutrino mixing framework the unitary PMNS neutrino matrix is parametrised by three angles and, depending on whether the massive neutrinos are Dirac of Majorana particles, by one Dirac or one Dirac and two Majorana CP-violation (CPV) phases. In the so-called “standard” parametrisation we will employ in the present study, the Pontecorvo–Maki–Nakagawa–Sakata (PMNS) matrix has the form [4, 3]:
@@
The approach to the lepton mixing problem based on non-Abelian discrete flavour symmetries combined with the gCP symmetry can lead to sharp predictions for lepton mixing angles and CPV phases. In this work, we shall confront the theoretical predictions of models based on this approach with the first JUNO results and the results of latest analysis of global neutrino data NuFIT-v6.0 [6], and we will perform statistical analysis of the predictions of the considered models. We shall focus on the breaking patterns of the non-Abelian discrete flavour and CP symmetries, which enforce the lepton mixing matrix to depend on only one or two real free parameters, concentrating on models with massive Majorana neutrinos. In these cases the six observables of the PMNS matrix, the three mixing angles and three CPV phases, are expressed in terms of one or two continuous real parameters and, in certain models, on one or two discrete parameters taking limited number of fixed values. As a result, some of the lepton mixing angles and CPV phases are strongly correlated with each other. These correlations can be tested experimentally. Besides the solar neutrino mixing angle $\theta_{12}$, we study the predictions for the atmospheric neutrino angle $\theta_{23}$, Dirac CPV phase $\delta_{CP}$ and Majorana CPV phases $\alpha_{21}$ and $\alpha_{31}$, as well as for the CPV rephasing invariant $J_{CP}$ which controls the magnitude of CPV effects in neutrino oscillations [45] and is a leptonic analog of the rephasing invariant introduced by Jarlskog in the quark sector [46]. It is interesting to observe that the JUNO’s first result can already strongly disfavor or even exclude some mixing patterns predicted by the breaking of the considered flavour and CP symmetries. For example, the lepton mixing matrix with the second column given by $(\frac{1}{\sqrt{3}},\frac{1}{\sqrt{3}},\frac{1}{\sqrt{3}})^{T}$ is ruled out by JUNO, although it is compatible with NuFIT-v6.0 at $3\sigma$ confidence level. Moreover, we see that the high precision measurement of $\theta_{12}$ by JUNO after six years of data collection, and the precise measurement of $\theta_{23}$ and $\delta_{CP}$ by the future long baseline experiments DUNE and T2HK allow to perform exhaustive test of the approach to lepton mixing based on the non-Abelian discrete flavour and CP symmetries. The JUNO’s result has been studied in the context of discrete flavour symmetry in the case of when a $Z_{2}$ residual symmetry is assumed to be preserved by the neutrino mass term [35, 42]. In this work, we shall perform a comprehensive analysis of possible symmetry breaking patterns in models in which a combination of non-Abelian discrete flavour symmetry and CP symmetry are assumed to hold initially. The potential of testing the non-Abelian discrete symmetry approach to the lepton flavour problem by the prospective data from JUNO, DUNE and T2HK has been investigated in [47], where only models based on the $A_{4}$ ($T^{\prime}$), $S_{4}$ and $A_{5}$ symmetries, but not also on the gCP symmetry, were considered. The capability of ESSnuSB to test discriminate between a class of lepton flavor models based on discrete non-Abelian flavor symmetrie was studied in [48]. In the models considered in the present study with discrete flavour and gCP symmetries, one obtains predictions for the Majorana phases $\alpha_{21}$ and $\alpha_{31}$ as well and sharper predictions for the Dirac phase $\delta_{CP}$.
@@
The structure of this paper is organised as follows. In section 2 the framework of determining lepton mixing from non-Abelian discrete flavour and CP symmetry breaking is presented, and we give the master formula of the lepton mixing matrix for different symmetry breaking patterns. The symmetry breaking pattern with $\mathcal{G}_{\ell}=Z_{n}\;(n\geq 3)$ or $K_{4}$, $\mathcal{G}_{\nu}=Z_{2}\times H^{\nu}_{CP}$ and the possible lepton mixing matrices are studied in section 3, and expressions of lepton mixing parameters as well as CP invariants for each case are presented. In this scenario, the lepton mixing matrix depends on a single real rotation angle $\theta$, and one column of the lepton mixing matrix is fixed by the residual symmetry. We study the symmetry breaking pattern $\mathcal{G}_{\ell}=Z_{2}\times H_{CP}^{\ell}$, $\mathcal{G}_{\nu}=Z_{2}\times H_{CP}^{\nu}$ in section 4, the resulting lepton mixing matrix depends on two free rotation angles $\theta_{\ell}$ and $\theta_{\nu}$, and only one entry of the lepton mixing matrix is fixed by symmetry consideration. The current neutrino oscillation data can be well accommodated in this scenario, and we make definite predictions for $\theta_{23}$ and $\delta_{CP}$. A statistical analysis is performed for both NO and IO neutrino mass spectra in section 5, the best fit values and the $3\sigma$ allowed regions of the mixing parameters are listed in tables for both neutrino mass spectra. The $\chi^{2}$ distributions of $\sin^{2}\theta_{12}$, $\sin^{2}\theta_{23}$ and $\delta_{CP}$ are also plotted. Finally we present our main conclusion in section 6. The details on the flavour groups $A_{5}$, $\Sigma(168)$ and $\Delta(6n^{2})$ are given in Appendix A. The allowed regions of the lepton mixing parameters for all the viable one-parameter and two-parameters models are collected in Appendix B. We have explored an alternative symmetry breaking pattern $\mathcal{G}_{\ell}=Z_{2}$, $\mathcal{G}_{\nu}=K_{4}\times H^{\nu}_{CP}$ in Appendix C, for which the resulting lepton mixing matrix also depends on two real free parameters. However, no phenomenologically viable lepton mixing matrix can be obtained in this scenario because the predicted value of the solar neutrino mixing angle $\theta_{12}$ lies significantly outside the $3\sigma$ region allowed by JUNO.
@@ 3 Lepton mixing with one free parameter
The possible residual subgroups $\mathcal{G}_{\ell}=Z_{n}$ or $K_{4}$ and $\mathcal{G}_{\nu}=Z_{2}\times H^{\nu}_{CP}$ have been studied for the symmetries $A_{5}\rtimes H_{CP}$, $\Sigma(168)\rtimes H_{CP}$ and $\Delta(6n^{2})\rtimes H_{CP}$ [60], and five possible patterns of lepton mixing matrix are found to be compatible with current experimental data of neutrino oscillation [60]. For phenomenological analysis in the next section, we report the explicit form of the mixing matrix together with the residual symmetry in the following. In addition, we shall give the analytical expressions for mixing angles and CP-violation (CPV) invariants $J_{CP}$, $I_{1}$ and $I_{2}$, associated with the Dirac and two Majorana CPV phases in the PMNS neutrino mixing matrix.
In the present work, we will adopt the standard parametrization of the lepton mixing matrix [3],
where $c_{ij}\equiv\cos\theta_{ij}$, $s_{ij}\equiv\sin\theta_{ij}$, $\delta_{CP}$ is the Dirac CPV phase, and $\alpha_{21,31}$ are Majorana CPV phases. For each independent mixing pattern, we can straightforwardly extract the expressions of the mixing angles $\sin^{2}\theta_{13}$, $\sin^{2}\theta_{12}$, $\sin^{2}\theta_{23}$ and the CP invariants $J_{CP}$, $I_{1}$, $I_{2}$ in the usual way. As regards the CPV, three weak basis invariants $J_{CP}$ [46], $I_{1}$ and $I_{2}$ [81, 82, 83, 84, 85] associated with the CP phases $\delta_{CP}$, $\alpha_{21}$ and $\alpha_{31}$ respectively can be defined
@@
Here the CPV rephasing invariant $J_{CP}$ which controls the magnitude of CPV effects in neutrino oscillations [45], is a leptonic analog of the rephasing invariant introduced by Jarlskog in the quark sector [46]. We note that a given Majorana phase $\alpha_{i1}$ would violate the CP-invariance if both ${\rm Im}(U^{*2}_{11}U^{2}_{1i})\neq 0$ and ${\rm Re}(U^{*2}_{11}U^{2}_{1i})\neq 0$, $i=2,3$ (see, e.g., [86]).
@@
The smallest symmetry group capable of reproducing the experimental mixing angles for specific values of $\theta$ is $S_{4}$. The second column of $U^{II}_{1}$ and $U^{II}_{2}$ are $(1,1,1)^{T}/\sqrt{3}$ which is the second column of tri-bimaximal mixing matrix, and consequently they are the so-called TM2 pattern. We can extract the following results for the lepton mixing angles
@@
This implies a lower bound on the solar mixing angle $\sin^{2}\theta_{12}>1/3$. Using the $3\sigma$ range of $\sin^{2}\theta_{13}$ from [6], we obtain $0.3402\leq\sin^{2}\theta_{12}\leq 0.3415$. The JUNO experiment has determined the $3\sigma$ range of the solar mixing angle $\sin^{2}\theta_{12}$ to be $[0.2831,,0.3353]$ [30], thus this mixing pattern has been ruled out by the JUNO experiment. Not only that, the TM2 mixing matrix has been excluded by the JUNO experiment [34, 39, 40]. Moreover, the reactor mixing angle and the atmospheric mixing angle are related as follow
@@
The flavour symmetry group $G_{f}=S_{4}$ yields the mixing patterns $U^{II}_{1}$ and $U^{II}_{2}$ for the parameter values $(\varphi_{1},\varphi_{2})=(\pi,0)$. In this configuration, the atmospheric mixing angle $\theta_{23}$ and the Dirac CPV phase $\delta_{CP}$ are both predicted to be maximal. Furthermore, the Majorana phases are constrained to be either 0 or $\pi$. Notably, $U^{II}_{1}$ and $U^{II}_{2}$ represent the same physical mixing pattern in this case, as they are connected through a redefinition of the parameter $\theta$ and the matrix $Q_{\nu}$:
@@
After some tedious calculations, we find the following relations between the mixing angles
@@
According to Eq. (3.41), the minimum value of the reactor mixing angle $\sin^{2}\theta_{13}$ is found to be 0.02153. Using the $3\sigma$ range of the reactor mixing angle $0.02153\leq\sin^{2}\theta_{13}\leq 0.02388$, we have
@@
$0.5405\leq\sin^{2}\theta_{23}\leq 0.5953~~\text{for}~U^{V}_{1}\,$,
$0.435\leq\sin^{2}\theta_{23}\leq 0.4595~~\text{for}~U^{V}_{2}\,$, (3.44)
Obviously the atmospheric mixing angle $\theta_{23}$ is non-maximal in this case. Note that The specific value of the solar mixing angle $\theta_{12}$ predicted by this mixing model has been ruled out by the latest results from the JUNO experiment [30].
We find the sum rules in Eq. (3.42) and consequently the estimates given in Eq. (3.44) are satisfied as well. Furthermore, the sum rule of Eq. (3.24) among the mixing angles and Dirac CP phase is fulfilled for all the four permutations of the PMNS matrices in Eq. (3.39). Consequently the comments below Eq. (3.24) also hold true here.

### v2

@@ Abstract
Working within the reference three-neutrino mixing framework, we confront the lepton mixing predictions derived using non-Abelian discrete flavour and CP symmetries with the first JUNO data on the solar neutrino mixing parameters $\sin^{2}\theta_{12}$ and with the results of the latest global neutrino data analysis. We focus on symmetry breaking patterns for which the lepton PMNS mixing matrix depends only on one or two free real parameters. Performing a comprehensive statistical analysis in each of the considered cases, we report the best fit values, the $3\sigma$ C.L. allowed ranges and the $\chi^{2}$-distributions of the lepton mixing observables - the three mixing angles and the three CP-violation phases. We find that the JUNO measurements can disfavour or rule out a number of the mixing patterns associated with specific types of breaking of the discrete flavour and CP symmetries. The synergy of JUNO, DUNE and T2HK data can provide an exhaustive test of the considered approach to lepton mixing based on non-Abelian discrete lepton flavour symmetries combined with the CP symmetry.
@@ 1 Introduction
The neutrino oscillation data accumulated over many years of measurements following the discovery of neutrino oscillations[1, 2, 3, 4, 5] revealed the existence of three-flavour neutrino mixing involving three light massive neutrinos (see, e.g., [6]). In this reference three-neutrino mixing framework the unitary PMNS neutrino matrix is parametrised by three angles and, depending on whether the massive neutrinos are Dirac of Majorana particles, by one Dirac or one Dirac and two Majorana CP-violation (CPV) phases. In the so-called “standard” parametrisation we will employ in the present study, the Pontecorvo–Maki–Nakagawa–Sakata (PMNS) matrix has the form [7, 6]:
@@
The approach to the lepton mixing problem based on non-Abelian discrete flavour symmetries combined with the gCP symmetry can lead to sharp predictions for lepton mixing angles and CPV phases. In this work, we shall confront the theoretical predictions of models based on this approach with the first JUNO results and the results of latest analysis of global neutrino data NuFIT-v6.0 [9], and we will perform statistical analysis of the predictions of the considered models. We shall focus on the breaking patterns of the non-Abelian discrete flavour and CP symmetries, which enforce the lepton mixing matrix to depend on only one or two real free parameters, concentrating on models with massive Majorana neutrinos. In these cases the six observables of the PMNS matrix, the three mixing angles and three CPV phases, are expressed in terms of one or two continuous real parameters and, in certain models, on one or two discrete parameters taking limited number of fixed values. As a result, some of the lepton mixing angles and CPV phases are strongly correlated with each other. These correlations can be tested experimentally. Besides the solar neutrino mixing angle $\theta_{12}$, we study the predictions for the atmospheric neutrino angle $\theta_{23}$, Dirac CPV phase $\delta_{CP}$ and Majorana CPV phases $\alpha_{21}$ and $\alpha_{31}$, as well as for the CPV rephasing invariant $J_{CP}$ which controls the magnitude of CPV effects in neutrino oscillations [48] and is a leptonic analog of the rephasing invariant introduced by Jarlskog in the quark sector [49]. It is interesting to observe that the JUNO’s first result can already strongly disfavor or even exclude some mixing patterns predicted by the breaking of the considered flavour and CP symmetries. For example, the lepton mixing matrix with the second column given by $(\frac{1}{\sqrt{3}},\frac{1}{\sqrt{3}},\frac{1}{\sqrt{3}})^{T}$ is ruled out by JUNO, although it is compatible with NuFIT-v6.0 at $3\sigma$ confidence level. Moreover, we see that the high precision measurement of $\theta_{12}$ by JUNO after six years of data collection, and the precise measurement of $\theta_{23}$ and $\delta_{CP}$ by the future long baseline experiments DUNE and T2HK allow to perform exhaustive test of the approach to lepton mixing based on the non-Abelian discrete flavour and CP symmetries. The JUNO’s result has been studied in the context of discrete flavour symmetry in the case of when a $Z_{2}$ residual symmetry is assumed to be preserved by the neutrino mass term [38, 45]. In this work, we shall perform a comprehensive analysis of possible symmetry breaking patterns in models in which a combination of non-Abelian discrete flavour symmetry and CP symmetry are assumed to hold initially. The potential of testing the non-Abelian discrete symmetry approach to the lepton flavour problem by the prospective data from JUNO, DUNE and T2HK has been investigated in [50], where only models based on the $A_{4}$ ($T^{\prime}$), $S_{4}$ and $A_{5}$ symmetries, but not also on the gCP symmetry, were considered. The capability of the ESSnuSB experiment to test and discriminate between a class of lepton flavor models based on discrete non-Abelian flavor symmetries was studied in [51]. In the models considered in the present study with discrete flavour and gCP symmetries, one obtains predictions for the Majorana phases $\alpha_{21}$ and $\alpha_{31}$ as well and sharper predictions for the Dirac phase $\delta_{CP}$.
@@
The structure of this paper is organised as follows. In section 2 the framework of determining lepton mixing from non-Abelian discrete flavour and CP symmetry breaking is presented, and we give the master formula of the lepton mixing matrix for different symmetry breaking patterns. The symmetry breaking pattern with $\mathcal{G}_{\ell}=Z_{n}\;(n\geq 3)$ or $K_{4}$, $\mathcal{G}_{\nu}=Z_{2}\times H^{\nu}_{CP}$ and the possible lepton mixing matrices are studied in section 3, and expressions of lepton mixing parameters as well as CP invariants for each case are presented. In this scenario, the lepton mixing matrix depends on a single real rotation angle $\theta$, and one column of the lepton mixing matrix is fixed by the residual symmetry. We study the symmetry breaking pattern $\mathcal{G}_{\ell}=Z_{2}\times H_{CP}^{\ell}$, $\mathcal{G}_{\nu}=Z_{2}\times H_{CP}^{\nu}$ in section 4, the resulting lepton mixing matrix depends on two free rotation angles $\theta_{\ell}$ and $\theta_{\nu}$, and only one entry of the lepton mixing matrix is fixed by symmetry consideration. The current neutrino oscillation data can be well accommodated in this scenario, and we make definite predictions for $\theta_{23}$ and $\delta_{CP}$. A statistical analysis is performed for both NO and IO neutrino mass spectra in section 5, the best fit values and the $3\sigma$ allowed regions of the mixing parameters are listed in tables for both neutrino mass spectra. The $\chi^{2}$ distributions of $\sin^{2}\theta_{12}$, $\sin^{2}\theta_{23}$ and $\delta_{CP}$ are also plotted. Finally we present our main conclusion in section 6. The details on the flavour groups $A_{5}$, $\Sigma(168)$, $\Delta(6n^{2})$ and $D_{n}$ are given in Appendix A. The allowed regions of the lepton mixing parameters for all the viable one-parameter and two-parameters models are collected in Appendix B. We have explored an alternative symmetry breaking pattern $\mathcal{G}_{\ell}=Z_{2}$, $\mathcal{G}_{\nu}=K_{4}\times H^{\nu}_{CP}$ in Appendix C, for which the resulting lepton mixing matrix also depends on two real free parameters. However, no phenomenologically viable lepton mixing matrix can be obtained in this scenario because the predicted value of the solar neutrino mixing angle $\theta_{12}$ lies significantly outside the $3\sigma$ region allowed by JUNO.
@@ 2.2 $\mathcal{G}_{\ell}=Z_{2}\times H^{\ell}_{CP}$ and $\mathcal{G}_{\nu}=Z_{2}\times H^{\nu}_{CP}$
In the paradigm of discrete flavor and CP symmetries, the CP violation arises from the mismatch between the residual CP symmetries of the charged lepton and neutrino sectors. If some common residual CP transformation is shared by the charged lepton and neutrino sectors, the CP symmetry would be preserved, the Dirac and Majorana CP violation phases would be equal to 0 or $\pi$ [27, 24]. On the other hand, if no common residual CP transformation is shared by the charged lepton and neutrino sectors, the CP symmetry is not conserved and the Majorana and/or Dirac phases take CP-violating values.
@@ 3 Lepton mixing with one free parameter
The possible residual subgroups $\mathcal{G}_{\ell}=Z_{n}$ or $K_{4}$ and $\mathcal{G}_{\nu}=Z_{2}\times H^{\nu}_{CP}$ have been studied for the symmetries $A_{5}\rtimes H_{CP}$, $\Sigma(168)\rtimes H_{CP}$ and $\Delta(6n^{2})\rtimes H_{CP}$ [63], and five possible patterns of lepton mixing matrix are found to be compatible with current experimental data of neutrino oscillation [63]. For phenomenological analysis in the next section, we report the explicit form of the mixing matrix together with the residual symmetry in the following. For each independent mixing pattern, we can straightforwardly extract the expressions of the mixing angles $\sin^{2}\theta_{13}$, $\sin^{2}\theta_{12}$, $\sin^{2}\theta_{23}$ and the CP invariants $J_{CP}$, $I_{1}$, $I_{2}$ in the usual way. In the present work, we will adopt the standard parametrization of the lepton mixing matrix in Eq. (1.1). As regards the CPV, three weak basis invariants $J_{CP}$, associated with the Dirac CPV phase $\delta_{CP}$, which is a leptonic analog of the invariant introduced by Jarlskog in the quark sector [49] and controls the magnitude of the CPV effects in neutrino oscillations  [48], and $I_{1}$ and $I_{2}$ related to the Majorana phases $\alpha_{21}$ and $\alpha_{31}$ [84, 85, 86, 87, 88] can be defined as:
@@
We note that a given Majorana phase $\alpha_{i1}$ would violate the CP-invariance if both ${\rm Im}(U^{*2}_{11}U^{2}_{1i})\neq 0$ and ${\rm Re}(U^{*2}_{11}U^{2}_{1i})\neq 0$, $i=2,3$ (see, e.g., [89]).
@@
The smallest symmetry group capable of reproducing the experimental mixing angles for specific values of $\theta$ is $S_{4}$ (i.e., $\Delta(24)\cong S_{4}$ with $(\varphi_{1},\varphi_{2})=(0,\pi)$). The second column of $U^{II}_{1}$ and $U^{II}_{2}$ are $(1,1,1)^{T}/\sqrt{3}$ which is the second column of tri-bimaximal mixing matrix, and consequently they are the so-called TM2 pattern. We can extract the following results for the lepton mixing angles
@@
This implies a lower bound on the solar mixing angle $\sin^{2}\theta_{12}>1/3$. Eliminating the rotation angle $\theta$, we find that the reactor mixing angle and the atmospheric mixing angle are related as follow,
@@
The flavour symmetry group $G_{f}=S_{4}$ yields the mixing patterns $U^{II}_{1}$ and $U^{II}_{2}$ for the parameter values $(\varphi_{1},\varphi_{2})=(0,\pi)$. In this configuration, the atmospheric mixing angle $\theta_{23}$ and the Dirac CPV phase $\delta_{CP}$ are both predicted to be maximal. Furthermore, the Majorana phases are constrained to be either 0 or $\pi$. Notably, $U^{II}_{1}$ and $U^{II}_{2}$ represent the same physical mixing pattern in this case, as they are connected through a redefinition of the parameter $\theta$ and the matrix $Q_{\nu}$:
@@
Finally, from the sum rule given in Eq. (3.19), using the $3\sigma$ range of $\sin^{2}\theta_{13}$ from [9], we obtain $0.3402\leq\sin^{2}\theta_{12}\leq 0.3415$. The JUNO experiment has determined the $3\sigma$ range of the solar mixing angle $\sin^{2}\theta_{12}$ to be $[0.2831,0.3353]$ [33]. Thus, the solar mixing angle is predicted to be larger than the $3\sigma$ upper limit of JUNO. Therefore the mixing patterns $U^{II}_{1}$ and $U^{II}_{2}$, as well as the TM2 mixing matrix, are excluded by the JUNO data [37, 42, 43, 45], although they are marginally compatible with neutrino global analysis NuFIT-v6.0 [9]. In a concrete model in which either the mixing pattern $U^{II}_{1}$ or $U^{II}_{2}$ is realized, the experimental data could possibly be accommodated if the solar mixing angle undergoes appropriate corrections induced by higher dimensional operators etc [24].
@@
After some tedious calculations, we find the following relations between the mixing angles  [^7] The sum rules relating $\sin^{2}\theta_{12}$ and $\sin^{2}\theta_{13}$ in Eqs. (3.28) and (3.34) were derived also in [56] in models with $A_{5}$ symmetry but without gCP symmetry, which were analysed phenomenologically in [50, 45]. These models lead to different predictions for $\delta_{CP}$ and $\sin^{2}\theta_{23}$, thus are experimentally distinguishable from the models considered in the present work.:
@@
From the expression of reactor angle in Eq. (3.40), we see that the minimum value of the reactor mixing angle fulfill $\sin^{2}\theta_{13}\simeq 0.02153$ when $\theta\simeq 0.5617\pi$. Requiring this mixing pattern to be compatible with the global analysis NuFIT [9] at $3\sigma$ level, we find that the allowed regions of the three mixing angles for normal ordering as follow,
$0.02153\leq\sin^{2}\theta_{13}\leq 0.02388\,$,
@@
$0.5405\leq\sin^{2}\theta_{23}\leq 0.5850,~~\text{for}~U^{V}_{1}\,$,
$0.4350\leq\sin^{2}\theta_{23}\leq 0.4595~~\text{for}~U^{V}_{2}\,$, (3.43)
Obviously the atmospheric mixing angle $\theta_{23}$ is non-maximal in this case. What is more important, the specific value of the solar mixing angle $\theta_{12}$ predicted by this mixing model has been ruled out by the first results from the JUNO experiment [33], although it is marginally compatible with NuFIT [9]. However, agreement with the experimental data could possibly be achieved in a specific model producing this mixing pattern at leading order, if $\theta_{12}$ acquires moderate corrections. As regards the leptonic CP violation phases, the sum rule of Eq. (3.23) among the mixing angles and Dirac CP phase is satisfied for all the four PMNS matrices  [^8] The sum rules for $\sin^{2}\theta_{12}$ and $\cos\delta_{CP}$, Eqs. (3.41) and (3.23), are present in models based on $A_{4}$, $S_{4}$ and $A_{5}$ symmetries but without gCP symmetry as well [56]. However, they lead to different predictions for $\sin^{2}\theta_{23}$ and thus for $\delta_{CP}$ [50, 45], and therefore are experimentally distinguishable from the models considered by us. $U^{V}_{1}$, $U^{V}_{2}$, $U^{V}_{3}$ and $U^{V}_{4}$.